OpenStudy (queenbb):

The graph shows the number of laps that Jan runs as her coach times her. What is the slope of the line and what does it mean in this situation? Select from the drop-down menus to correctly complete each statement. The slope of the line is . This means that Jan runs laps every .

OpenStudy (blazeryder):

Where is the graph?

OpenStudy (queenbb):

one second

563blackghost (563blackghost):

\(\huge\bf{\frac{rise}{run} \rightarrow \frac{y_{2}-y_{1}}{x_{2}-x_{1}} = \color{red}{slope}}\)

OpenStudy (queenbb):

@RhondaSommer

563blackghost (563blackghost):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @563blackghost \(\huge\bf{\frac{rise}{run} \rightarrow \frac{y_{2}-y_{1}}{x_{2}-x_{1}} = \color{red}{slope}}\) \(\color{#0cbb34}{\text{End of Quote}}\) Use the process I typed out.

OpenStudy (queenbb):

help plz

563blackghost (563blackghost):

Plug in the points.

OpenStudy (queenbb):

what is the answer i am checkining for what i worte

563blackghost (563blackghost):

I'm sorry but I can not give direct answers since it is in violation of the Code of Conduct.

563blackghost (563blackghost):

I can only aid you.

OpenStudy (queenbb):

i need help @retirEEd

OpenStudy (queenbb):

@563blackghost i dont k now tto do that rise tun thing

563blackghost (563blackghost):

Just plug the points in... \(\huge\bf{(x_{1},y_{1}) \rightarrow (2,4)}\) \(\huge\bf{(x_{2},y_{2}) \rightarrow (5,10)}\) Plug in. \(\huge\bf{\frac{rise}{run} \rightarrow \frac{y_{2}-y_{1}}{x_{2}-x_{1}} \rightarrow \frac{10-4}{5-2}} = \color{red}{slope}\)